OpenStudy (anonymous):

if 2^x -3x +1 = a(x-1)^2 + b(x-1)+c find the values of a, b and c

OpenStudy (chaise):

Are you sure it is supposed to be 2^x?

OpenStudy (anonymous):

oopss my apologies, its meant to be 2x^2

OpenStudy (anonymous):

Expand right side of the equation, then equate coefficients of both expressions. You ok with this?

OpenStudy (anonymous):

@ stroodle, would you be kind enough to show me the working out?

OpenStudy (chaise):

If: 2x^2-3x+1=a(x-1)^2 + b(x-1)+c Make both equations equal to 0, when you do that they will be both equal to eachother. 2x^2-3x+1=0 a(x-1)^2+b(x-1)+c=0 Now you can solve the first equation. x = 1 or x=1/2 Now you can substitute your x into that equation.. I'm not sure what to do from here.. but hopefully this helps. I might be wrong though.

OpenStudy (anonymous):

brb. I'll post another solution in 10-15mins if no-one else does.

OpenStudy (chaise):

You have four unknowns, you will need to somehow create for individual equations from that equation and solve them simultaneously. Possibly.

OpenStudy (anonymous):

umm thanks chaise. lets wait and see how stroodle does it! stroodle your input is highly anticipated!

OpenStudy (anonymous):

if you expand rearrange the right side of the equation you should get: \[ax^2+(b-2a)x+a+c-b\] So now you can equate the coefficients of the two sides of the equation. i.e. a=2 b-2a=-3 a+c-b=1

OpenStudy (anonymous):

Let me know if you don't get it.

OpenStudy (anonymous):

why thank you @ stroodle i fully understand, thank you math whiz!